OpenStudy (anonymous):

In quadratic functions: f(x) = x^2 does 2f(x) = f(-2x)?

OpenStudy (mertsj):

no

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

correction does 2f(x) = f(2x) when f(x) = x^2

OpenStudy (mertsj):

no

OpenStudy (anonymous):

still no

OpenStudy (anonymous):

ok, for what x values is it different?

OpenStudy (mertsj):

Try 3

OpenStudy (anonymous):

2 f(x) would be 2x^2 f(2x) would be 4x^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i'm evaluating it incorrectly then

OpenStudy (anonymous):

i tried 3 and got the same answers

OpenStudy (anonymous):

2f(x), when x=3 2(3^2) f(2x), when x=3 f(2*3^2)

OpenStudy (mertsj):

f(x)=x^2 f(3)=3^2=9 2f(3)=2(9)=18 f(2x)=f(6)=6^2=36 18 does not equal 36

OpenStudy (anonymous):

2f(x), when x=3 2(3^2) 2(9) 18 f(2x), when x=3 f(2*3^2) (2*9) 18

OpenStudy (anonymous):

ok ty

OpenStudy (mertsj):

yw